solve by substituting. x^-2-6x^-1=-4. hint: u=x^-1
Answer by jim_thompson5910(35256)

u=x^-1

u^2=x^-2

-------------------------------------------------------

So

x^-2-6x^-1=-4

becomes

u^2-6u = -4

which turns into

u^2 - 6u + 4 = 0

Now use the quadratic formula to solve for 'u'. Once you have two solutions for 'u', use them to find the two solutions for 'x'.